y" - y = e2x[3 tan e+ 3 (sec e*)2]


1
Expert's answer
2021-03-30T08:49:08-0400

"\displaystyle
y'' - y = e^{2x}\left(3\tan{e}+ \frac{3(\sec{e})}{2}\right) = Ce^{2x} \\

m^2 - 1 = 0 \\

(m + 1)(m - 1) = 0\\

m = -1,\,\, 1\\

\textsf{The complementary factor is}\,\,\, y = A\cosh{x} + B\sinh{x} \\

\textsf{The particular integral is}\,\,\, y = De^{2x} \\/

y'' = 4De^{2x},\,\, y = De^{2x} \\

4De^{2x} - De^{2x} = Ce^{2x} \\

3De^{2x} = Ce^{2x} \\

D = \frac{C}{3}\\

y = \frac{e^{2x}\left(3\tan{e}+ \frac{3(\sec{e})}{2}\right)}{3} \\

\therefore y = C.F + P.I \\

y = A\cosh{x} + B\sinh{x} + \frac{e^{2x}\left(3\tan{e}+ \frac{3(\sec{e})}{2}\right)}{3}.\\"